OpenStudy (anonymous):

A test of the prototype of a new automobile shows that the minimum distance for a controlled stop from 100 km/h to zero is 41 m. Find the acceleration, assuming it to be constant as a fraction of the free-fall acceleration. The acceleration due to gravity is 9.81 m/s 2 . Answer in units of g

OpenStudy (anonymous):

Average acceleration is \(\overline a = \dfrac{\Delta v}{\Delta t}\). Then, divide out \(g\) to get in units of \(g\).

OpenStudy (anonymous):

Ok, my numbers are 93 km/h and 50 m to zero. I tried to follow and this is what I got which is wrong... 50m=93t, t= 0.537634m(60s)= 32.258065 s a= (0-93)/(32.258065s) = -2.882999 for final answer that was wrong Also the system wants it rounded to 6 decimal places... What did I do wrong?